# Non-linear shattered pellet injection modelling in ASDEX Upgrade

### v1

@@ 3.1 Simulation set-up
In order to prevent extremely large $\eta$ values, a temperature threshold $T_{min}$ = 1 eV is used to cut off the temperature dependency, i.e., when the temperature drops below this value, $\eta$ will not further increase. In the same vein, an upper limit of temperature dependency is applied at $T_{max}$ = 1.35 keV, corresponding to a $\eta$ floor value of $2~{}\times~{}10^{-8}~{}\Omega$m to simplify the simulation numerically. The perpendicular viscosity $\mu$ and parallel viscosity $\mu_{\parallel}$ are both set as $4.85\times 10^{-7}~{}\mathrm{kg~{}m^{-1}s^{-1}}$. The corresponding kinematic viscosity in the center $\nu_{0}=\mu/\rho_{0}$ is $1.6\ \mathrm{m^{2}/s}$. The perpendicular thermal diffusivity is set as $\chi_{\perp}=1\ \mathrm{m^{2}/s}$. Spitzer-Haerm parallel thermal diffusivity,
@@
is adopted in the parallel direction. A floor value at 50 eV is applied in the simulation to help deal with low temperature regions such as temperature holes. The isotropic particle diffusivity for main ions $D$ is set to be 1.6 $\mathrm{m^{2}/s}$. The impurity diffusivity is set to be the same with the main ions, i.e. $D_{imp}=1.6\ \mathrm{m^{2}/s}$. No background impurity is included in all the simulations to keep the focus on the injected material.
The simulation utilizes a flux-aligned grid with a resolution of 70 $\times$ 110 in the radial and poloidal direction, respectively. 8 toroidal harmonics n = 0…7 are included in our simulations. Since the simulations are aimed for late TQ to early CQ, an ideal conducting wall is assumed.
The SPI is launched at the very beginning in our simulations. Once a shard enters the plasma domain, the ablation rate for a mixed neon and deuterium shard is calculated based on an improved version [35] of the neutral gas shielding (NGS) model [36, 37]. The ablation rate of each shard is obtained by
@@
$\ \left(\frac{r_{p}[\mathrm{cm}]}{0.2}\right)^{4/3}n_{e}[10^{14}\mathrm{cm}^{-3}]^{1/3},$
where $X=N_{D_{2}}/(N_{D_{2}}+N_{Ne})$ is the molecular mixture ratio of deuterium and $r_{p}$ is the radius of the shard. The ablated material is assumed to be a Gaussian shape source term centered at the shard location, with the following form:
@@
In equation (12), $R_{\mathrm{s}},Z_{\mathrm{s}}$ and $\phi_{\mathrm{s}}$ are the real-time location of one specific shard. $\Delta r_{\mathrm{NG}}$ is the radius of the neutral gas cloud in the poloidal plane, which is set to 1.4×$\times r_{p}$ with a minimum value of 0.03 m, where $r_{p}$ is the radius of the shard. $\Delta\phi_{\mathrm{NG}}$ is the toroidal extent of the cloud, which is set to 0.8 rad in our simulation.
@@ 3.2 Neon content scan
To investigate the effect of the neon concentration, a scan in the neon content is performed at fixed number of fragments (LF_FV_Ne0, LF_FV_Ne0.12, LF_FV_Ne1 and LF_FV_Ne10 in Tab. 1). To isolate the effects of the different neon fractions, we assume an identical size and velocity of the shards in the scan. The total injected neon and deuterium atoms are listed in Tab. 1.
@@
Figure 3 shows the evolution of global quantities over time after the injection. From the thermal energy evolution in figure 3 (a), it can be seen that a very rapid drop of thermal energy is triggered upon the shards arriving at the plasma domain at about 0.45 ms, regardless of the neon fraction. The sudden drop of about 50% of the thermal energy is mainly due to the convective and conductive transport caused by the magnetic field line stochastization at the edge. For the pure deuterium case (0% neon), the plasma enters a prolonged non-disruptive phase following the initial drop in thermal energy. In contrast, for neon-doped pellet cases, the thermal energy is gradually depleted after the initial drop. The pre-TQ+TQ duration (period from shards arrival to the end of TQ) increases as the neon fraction decreases. The time of shard arrival and exit are marked by the purple and cyan vertical dashed lines, respectively. The time at the end of TQ is annotated by the black vertical dashed lines in the plot when the total thermal energy drops to about 100 kJ, which is in line with the the radiation peaks in figure 3 (e).
@@
In figure 3 (c), the total assimilated neon quantity is plotted, while the assimilation ratio of impurity neon, calculated by the ratio of assimilated neon to the total injected neon content, is shown in figure 3 (d). The 10% neon pellet has the largest total assimilated neon quantity, whereas its assimilation ratio is the lowest comparing to other neon fraction cases. This is a self-regulating process that, with a lower neon fraction, the electron temperature decreases more gradually, resulting in an extended ablation period and consequently a higher assimilation ratio, which is consistent with the INDEX results in Ref. [17]. It should be noted that for the 10% and 1% neon cases, ablation stops at approximately 2 and 3 ms, respectively, when the total thermal energy drops to nearly 100 kJ, in line with the black vertical dashed lines annotating the end of TQ. However, for the 0.12% neon case, ablation saturates at around 3.5 ms as marked by the cyan vertical dashed line, while the plasma temperature remains sufficiently high for shard ablation. This occurs because, at that moment, the majority of the shards have traversed the plasma and start to exit the plasma domain. It is worth noting that even for a pellet with a low neon fraction of 0.12%, the ablation ratio remains limited to approximately 20%, indicating that a significant amount of material still remains unablated. Further increasing the neon fraction to 20% would not enhance the cooling effect of the pellet much, due to the limitation imposed by ablation [38].
@@
Figure 3 (e) shows the evolution of radiation power versus time, where double radiation peaks are observed for the higher neon fraction cases (1% and 10%). Analysis of the radiation distribution reveals that the first radiation peak originates from the edge near the injection location, while the second peak is associated with the plasma core. The first radiation peak is due to the initial ablation of the shards, while the second peak is related to the 1/1 MHD mode during core collapse.
@@
To understand the two-stage cooling observed in our simulations, figure 4 displays the evolution of total radiated energy, ohmic heating energy, total thermal energy and radiation fraction for the cases with 0.12%, 1% and 10% neon. It shows that, before 1 ms (indicated by the gray dashed line), the radiation fraction is very limited. In the cases with 0.12%, 1%, and 10% neon, the radiation fractions at that time are approximately 0%, 6%, and 27%, respectively. During the second stage of thermal energy drop (after 1 ms), energy loss is dominated by radiative cooling, with the radiation fraction increasing from 80% to 97% as the neon content increases. Here, the radiation fraction is calculated using the formula $\Delta W_{rad}/(\Delta W_{th}+\Delta W_{ohm})$, where $\Delta W_{rad}$ represents the change in radiated energy, $\Delta W_{th}$ is the change in thermal energy, and $\Delta W_{ohm}$ refers to the change in Ohmic heating energy. The calculation is separated into two distinct time windows of the thermal energy drop. Due to the absence of background impurity and lack of impurity radiation, the second phase may not occur at all in pure deuterium injection, like also observed in many non-disruptive shots in experiments [20].
@@
To study the first stage of thermal energy loss in more depth, the Poincaré plots of magnetic fields for the pure deuterium injection case are plotted in figure 5. It can be seen that right after the vanguard shards reach the plasma edge (0.48 - 0.52 ms), the magnetic field becomes stochastic near the shard locations, due to the destabilization by helical cooling. Since the injection targets a H-mode plasma, the large current density gradient and pressure gradient in the pedestal are prone to MHD instabilities, triggering magnetic reconnection and leading to the stochastization. Due to the steep pressure gradient at the pedestal, consequently, considerable energy loss occurs during the early injection phase once confinement at edge is broken. Afterwards, when shards reach the q = 2 surface at around t = 0.57 ms, a large 2/1 island is induced and triggers the global reconnection event, resulting in full stochastisation of the flux surface. This large 2/1 island can be seen by producing Poincaré plots, which include only the n=0 and n=1 toroidal harmonics of the magnetic field, as shown in figure 6 (b). In figure 6 (a), the perturbed poloidal magnetic flux $\delta\psi$ at the same time is plotted, revealing a 2/1 tearing mode structure.
@@
The evolution of electron temperature and density during first stage of thermal energy loss are shown in figure 7, and the flux averaged electron temperature profiles are plotted in figure 8. Upon the vanguard shards arriving at the q = 3 surface at 0.479 ms, the 3/1 magnetic islands are induced, as the m = 3 structure of the ablated electron density becomes clearly visible in figure 7 (a). The pedestal is subsequently degraded, and the temperature becomes flattened near the q = 3 resonant surface at around t = 0.5 ms, as shown in figure 7 (b) and in the flux averaged profile in figure 8. Following the pedestal collapse, significant heat flux propagates outward to the edge, causing a sudden ablation of the subsequent shards. This is illustrated in figure 7(c) and (d), where considerable thermal energy is expelled from the edge. Figure 8 also indicates that the majority of thermal energy loss during the first stage occurs between 0.5 and 0.6 ms. The evolution of the temperature profile is similar across different neon fractions, due to the absence of significant radiative cooling at this stage.
@@
During the first stage of the cooling, the majority of the shards are located at the plasma edge. Despite significant degradation, the plasma temperature is still warm enough ($\sim$ 500 eV) for ablation to continue. With impurity radiation coming into effect, the second stage of cooling diverges depending on the neon fraction. The magnetic energy spectrum is compared in figure 9, for the four cases. In the pure deuterium case, only a single MHD peak is observed during the initial injection stage, corresponding to the rapid drop in thermal energy at the beginning. For neon-doped injections, in addition to the first MHD peak observed at the early stage, a later MHD peak appears at the end of the TQ phase, aligning with the radiation peaks. The second MHD peak corresponds to a 1/1 mode activity, which leads to the core collapse. In cases with a higher neon fraction, when the shards reach the core, the core collapse is triggered shortly due to stronger MHD response. With low neon concentration, the radiative cooling is less intense, leading to milder MHD activity. As a result, the plasma remains more stable for a longer period, and the core collapse occurs after the shards have exited the plasma domain.
@@
This slow TQ process can also be seen in figure 10, where the temporal evolution of the flux averaged temperature profiles are compared for different neon fractions. For all these cases, after the initial degradation of temperature, an inside-out cooling pattern is observed. As the shards penetrate inward, they create a localized low-temperature region that moves inward along with the shards, while a temperature layer of approximately 100 eV remains at the edge. This inside-out cooling does not occur at higher shard numbers or lower penetration speed, as discussed later. For pure deuterium injection, due to lack of radiation, the thermal energy is maintained and the temperature profile gradually relaxes over time. In contrast, with neon mixed injections, the core temperature suddenly drops to below 10 eV following the 1/1 MHD activity, and then initiating the inside-out cooling.
3.3 Influence of shard size
@@
To understand the influence of shard size on the disruption mitigation effectiveness, simulations with different shard sizes are conducted with a neon fraction of 0.12% and 10%, corresponding to the cases MF_FV_Ne0.12, SF_FV_Ne0.12, MF_FV_Ne10 and SF_FV_Ne10 from Tab. 1.
@@
Figure 11: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity and (d) radiation power for different shard numbers for fixed total volume/material with 0.12% neon. The time of shard arrival $t_{arr}$ and exit $t_{ext}$ at/from the plasma domain are marked by the purple and cyan vertical dashed lines, respectively. The time at the end of TQ is annotated by the black vertical dashed lines.
@@
Figure 12: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity and (d) radiation power for different shard numbers for fixed total volume/material with 10% neon. The time of shard arrival $t_{arr}$ and exit $t_{ext}$ at/from the plasma domain are marked by the purple and cyan vertical dashed lines, respectively. The time at the end of TQ is annotated by the black vertical dashed lines.
@@
Figure 13: Flux averaged temperature profile evolution for different shard numbers with 0.12% neon. The dashed lines indicate the estimated shard trajectory.
@@
Figure 14: Flux averaged temperature profile evolution for different shard numbers with 10% neon. The dashed lines indicate the estimated shard trajectory.
@@
Figure 11 shows the overview plots for different shard numbers in the 0.12% neon scenario, where the total injected material is identical across cases. In figure 11 (a), a trend is observed that a larger number of smaller shards can lead to a shorter TQ duration as indicated by the thermal energy evolution. This agrees with the lower dimensional results we previously obtained between JOREK and INDEX [16]. The larger total surface area for smaller shards contributes to stronger assimilation until the majority of the shards exit the plasma at around 3.5 ms. Since the end of ablation is not determined by a full thermal collapse here, but by the leaving of the shards, the ablation for all cases saturate at almost the same time, which can be clearly seen in figure 11 (c). As the ablation duration is not constrained by the low temperature at TQ, smaller shards achieve higher assimilation compared to larger ones. Owing to the larger assimilation, the core collapse occurs earlier with smaller shards and stronger radiation power is observed at the core collapse phase, as shown in figure 11 (d).
@@
Our numerical results deviate from the experimental findings reported in [39], where larger shards with low neon fraction exhibit better material assimilation. The key reason for this discrepancy is that our current modelling fails to accurately simulate the plasmoid drift during the pre-TQ phase and dose not take into account the rocket effect, which accelerates the shard and deflects its trajectory. Smaller shards can drive stronger plasmoid drift due to quicker ablation and larger pressure peaking, and are more susceptible to the rocket effect due to smaller mass. In experiments [39], smaller shards tend to have longer pre-TQ duration yet achieve lower assimilation. Simulations using the NGS model indicate that smaller shards ablate at a higher rate; if their total assimilation remains lower despite longer ablation durations, it is very likely due to the plasmoid drift and the rocket effect expelling the material from the plasma for low neon fraction injections.
@@
Figure 12 shows the same data, but for the 10% neon cases with different shard numbers. It can be observed that a larger number of smaller shards leads to quicker assimilation due to the larger total surface area at the beginning, similar to cases with lower neon content. However, a reduced number of larger shards has a longer ablation duration, resulting in a relatively similar total assimilation level when ablation ceases. In contrast to the 0.12% neon cases, the ablation stops due to the temperature drop during the TQ, which occurs before the shards leave the plasma on the high field side. By the end of the TQ phase, the 199 shards case shows the highest assimilation, followed by the 1105 shards case, with the 53 shards case having the lowest assimilation. However, the trend is now less pronounced, and the assimilation levels are not significantly different, aligning with experimental observations that the size effect on assimilation is relatively small for pellets with higher neon fractions [39, 20]. This reaffirms our previous hypothesis, since the plasmoid drift is suppressed at high neon fractions. It should be noted that, in those 10% neon cases, a brief decline in assimilation is observed after the TQ. This is due to a transient increase in impurity diffusivity to 4 - 6 times its original value, i.e. $6.4-9.6\ \mathrm{m^{2}/s}$, to overcome numerical difficulties. Once this challenging period passes, the impurity diffusivity is restored to its original value. For the 199 shards case with 10% neon, the simulation could not proceed beyond the onset of the core collapse due to numerical issues.
@@
The temperature profile evolution for the 0.12% neon cases is shown in figure 13, and for the 10% neon cases in figure 14. The 10% neon case with 199 shards is excluded from the comparison, as this simulation did not complete the TQ. In the low neon fraction cases, an inside-out cooling pattern is observed across all configurations, owing to limited impurity assimilation at the edge. However, in the high neon fraction cases, as the number of shards increases, the cooling pattern transitions from inside-out to outside-in. The plume generated by the smaller shards is more elongated, ensuring a continuous presence of shards at the edge. Additionally, smaller shards exhibit more efficient assimilation during the initial stage. Consequently, a cold front moving alongside the pellets inwards is established more effectively with smaller shards.

### v2

@@ 3.1 Simulation setup
In order to prevent infinite $\eta$ values at extremely low temperature, a temperature threshold $T_{min}$ = 1 eV is used to cut off the temperature dependency, i.e., when the temperature drops below this value, $\eta$ will not further increase. In the same vein, an upper limit of temperature dependency is applied at $T_{max}$ = 1.35 keV, corresponding to a $\eta$ floor value of $2~{}\times~{}10^{-8}~{}\Omega$m to simplify the simulation numerically. The perpendicular viscosity $\mu$ and parallel viscosity $\mu_{\parallel}$ are both set as $4.85\times 10^{-7}~{}\mathrm{kg~{}m^{-1}s^{-1}}$. The corresponding kinematic viscosity in the center $\nu_{0}=\mu/\rho_{0}$ is $1.6\ \mathrm{m^{2}/s}$. The perpendicular thermal diffusivity is set as $\chi_{\perp}=1\ \mathrm{m^{2}/s}$. Spitzer-Haerm parallel thermal diffusivity,
@@
is adopted in the parallel direction. A floor value at 50 eV is applied in the simulation to help deal with low temperature regions such as temperature holes. The isotropic particle diffusivity for main ions $D$ is set to be 1.6 $\mathrm{m^{2}/s}$. The impurity diffusivity is set to be the same with the main ions, i.e. $D_{imp}=1.6\ \mathrm{m^{2}/s}$. No background impurity and no initial magnetic perturbation are included in the simulations to keep the focus on the injected material.
The simulations utilize a flux-aligned grid, as shown in figure 4, with a resolution of 70 $\times$ 110 in the radial and poloidal direction, respectively. After rigorous convergence tests on number of toroidal Fourier modes, 8 toroidal harmonics (n = 0…7) are chosen for the production simulations. Since the simulations are aimed for late TQ to early CQ, an ideal conducting wall is assumed. The nominal time step used in our simulations is approximately $6.1\times 10^{-8}$ s, which corresponds to 0.1 times the normalized JOREK time unit $t_{norm}$, during the CQ phase, the time step can be gradually increased to 1 - 2 times $t_{norm}$.
Before launching our production simulations, we conducted a preliminary assessment to ensure that the diffusive evolution of the plasma equilibrium would not interfere with the results. Specifically, we examined the evolution of the equilibrium profiles over a period of 1.2 ms without any injection. During this interval, the equilibrium changes only slightly, with no significant alteration in the global profiles. The SPI is launched at the very beginning in our simulations. Once a fragment enters the plasma domain, the ablation rate for a mixed neon and deuterium fragment is calculated based on an improved version [38] of the neutral gas shielding (NGS) model [39, 40]. The ablation rate of each fragment is obtained by
@@
×$\times\left(\frac{r_{p}[\mathrm{cm}]}{0.2}\right)^{4/3}\left(n_{e}[10^{14}\mathrm{cm}^{-3}]\right)^{1/3},$
where $X=N_{D_{2}}/(N_{D_{2}}+N_{Ne})$ is the molecular mixture ratio of deuterium and $r_{p}$ is the radius of the fragment. The ablated material is assumed to be a Gaussian shape source term centered at the fragment location, with the following form:
@@
In equation (12), $R_{\mathrm{s}},Z_{\mathrm{s}}$ and $\phi_{\mathrm{s}}$ are the real-time location of one specific fragment. $\Delta r_{\mathrm{NG}}$ is the radius of the neutral gas cloud in the poloidal plane, which is set to 0.03 m. The poloidal grid spacing in our setup is approximately 0.03 m, which already represents a very fine mesh given the global scale of the simulations. Using a narrower deposition width would not be numerically meaningful in this context, as it would fall below the grid resolution and lead to under-resolved or numerically unstable behavior. $\Delta\phi_{\mathrm{NG}}$ is the toroidal extent of the cloud, which is set to 0.8 rad in our simulation, due to the limitation of the toroidal grid resolution.
@@ 3.2 Neon content scan
To investigate the effect of the neon concentration, a scan in the neon content is performed at fixed number of fragments (LF_FV_Ne0, LF_FV_Ne0.12, LF_FV_Ne1 and LF_FV_Ne10 in Tab. 1). To isolate the effects of the different neon fractions, we assume an identical size and velocity of the fragments in the scan. The total injected neon and deuterium atoms are listed in Tab. 1.
@@
Figure 5 shows the evolution of global quantities over time after the injection. From the thermal energy evolution in figure 5 (a), it can be seen that a very rapid drop of thermal energy is triggered upon the fragments arriving at the plasma domain at about 0.45 ms, regardless of the neon fraction. The sudden drop of about 50% of the thermal energy is mainly due to the convective and conductive transport caused by the magnetic field line stochastization at the edge. For the pure deuterium case (0% neon), the plasma enters a prolonged non-disruptive phase following the initial drop in thermal energy. In contrast, for neon-doped pellet cases, the thermal energy is gradually depleted after the initial drop. The pre-TQ+TQ duration (period from fragments arrival to the end of TQ) increases as the neon fraction decreases. The time of fragment arrival and exit are marked by the purple and cyan vertical dashed lines, respectively. The times at the end of TQ are annotated by the colorful vertical dashed lines in the plot when the total thermal energy drops to about 100 kJ, which are in line with the the radiation peaks in figure 5 (e).
@@
In figure 5 (c), the total assimilated neon quantity is plotted, while the assimilation ratio of impurity neon, calculated by the ratio of assimilated neon to the total injected neon content, is shown in figure 5 (d). The 10% neon pellet has the largest total assimilated neon quantity, whereas its assimilation ratio is the lowest comparing to other neon fraction cases. This is a self-regulating process that, with a lower neon fraction, the electron temperature decreases more gradually, resulting in an extended ablation period and consequently a higher assimilation ratio, which is consistent with the INDEX results in Ref. [17]. It should be noted that for the 10% and 1% neon cases, ablation stops at approximately 2 and 3 ms, respectively, when the total thermal energy drops to nearly 100 kJ at the end of TQ. However, for the 0.12% neon case, ablation saturates at around 3.5 ms as marked by the cyan vertical dashed line, while the plasma temperature remains sufficiently high for fragment ablation. This occurs because, at that moment, the majority of the fragments have traversed the plasma and start to exit the plasma domain. It is worth noting that even for a pellet with a low neon fraction of 0.12%, the ablation ratio remains limited to approximately 20%, indicating that a significant amount of material still remains unablated in the simulation.
@@
Figure 5 (e) shows the evolution of radiation power versus time, where double radiation peaks are observed for the higher neon fraction cases (1% and 10%). Analysis of the radiation distribution reveals that the first radiation peak originates from the edge near the injection location, while the second peak is associated with the plasma core. The first radiation peak is due to the initial ablation of the fragments, while the second peak is related to the 1/1 MHD mode during core collapse.
@@
To understand the two-stage cooling observed in our simulations, figure 6 displays the evolution of total radiated energy, ohmic heating energy, total thermal energy and radiation fraction for the cases with 0.12%, 1% and 10% neon. It shows that, before 1 ms (indicated by the gray dashed line), the radiation fraction is very limited. In the cases with 0.12%, 1%, and 10% neon, the radiation fractions at that time are approximately 0%, 6%, and 27%, respectively. During the second stage of thermal energy drop (after 1 ms), energy loss is dominated by radiative cooling, with the radiation fraction increasing to 80%, 93% and 97%, respectively, as the neon content increases. Here, the radiation fraction is calculated using the formula $\Delta W_{rad}/(\Delta W_{th}+\Delta W_{ohm})$, where $\Delta W_{rad}$ represents the change in radiated energy, $\Delta W_{th}$ is the change in thermal energy, and $\Delta W_{ohm}$ refers to the change in Ohmic heating energy. The calculation is separated into two distinct time windows of the thermal energy drop. Due to the absence of background impurity and lack of impurity radiation, the second phase may not occur at all in pure deuterium injection, like also observed in many non-disruptive shots in experiments [20].
@@
To study the first stage of thermal energy loss in more depth, the Poincaré plots of magnetic fields for the pure deuterium injection case are plotted in figure 7. The plot shows that right after the vanguard fragments reach the plasma edge (0.48 - 0.52 ms), the magnetic field becomes stochastic near the fragment locations, due to the destabilization by helical cooling. Since the injection targets a H-mode plasma, the large current density gradient and pressure gradient in the pedestal are prone to MHD instabilities, triggering magnetic reconnection and leading to the stochastization. Due to the steep pressure gradient at the pedestal, consequently, considerable energy loss occurs during the early injection phase once confinement at edge is broken. Afterwards, when fragments reach the q = 2 surface at around t = 0.57 ms, a large 2/1 island is induced and triggers the global reconnection event, resulting in full stochastisation of the flux surface. This large 2/1 island can be seen by producing Poincaré plots, which include only the n=0 and n=1 toroidal harmonics of the magnetic field, as shown in figure 8 (b). In figure 8 (a), the perturbed poloidal magnetic flux $\delta\psi$ at the same time is plotted, revealing a 2/1 tearing mode structure.
@@
The evolution of electron temperature and density the during first stage of thermal energy loss are shown in figure 9, and the flux-averaged electron temperature profiles are plotted in figure 10. Upon the vanguard fragments arriving at the q = 3 surface at 0.479 ms, the 3/1 magnetic islands are induced, as the m = 3 structure of the ablated electron density becomes clearly visible in figure 9 (a). The pedestal is subsequently degraded, and the temperature becomes flattened near the q = 3 resonant surface at around t = 0.5 ms, as shown in figure 9 (b) and in the flux-averaged profile in figure 10. Following the pedestal collapse, significant heat flux propagates outward to the edge, causing a sudden ablation of the subsequent fragments. This is illustrated in figure 9(c) and (d), where considerable thermal energy is expelled from the edge. Figure 10 also indicates that the majority of thermal energy loss during the first stage occurs between 0.5 and 0.6 ms. The evolution of the temperature profile is similar across different neon fractions, due to the absence of significant radiative cooling at this stage.
@@
During the first stage of the cooling, the majority of the fragments are located at the plasma edge. Despite significant degradation, the plasma temperature is still warm enough ($\approx$ 500 eV) for ablation to continue. With impurity radiation coming into effect, the second stage of cooling diverges depending on the neon fraction. The magnetic energy spectrum is compared in figure 11, for the four cases. In the pure deuterium case, only a single MHD peak is observed during the initial injection stage, corresponding to the rapid drop in thermal energy at the beginning. For neon-doped injections, in addition to the first MHD peak observed at the early stage, a later MHD peak appears at the end of the TQ phase, aligning with the radiation peaks. The second MHD peak corresponds to a 1/1 mode activity, which leads to the core collapse. In cases with a higher neon fraction, when the fragments reach the core, the core collapse is triggered shortly due to stronger MHD response. With low neon concentration, the radiative cooling is less intense, leading to milder MHD activity. As a result, the plasma remains more stable for a longer period, and the core collapse occurs after the fragments have exited the plasma domain.
@@
This slow TQ process can also be seen in figure 12, where the temporal evolution of the flux-averaged temperature profiles are compared for different neon fractions. For all these cases, after the initial degradation of temperature, an inside-out cooling pattern is observed. As the fragments penetrate inward, they create a localized low-temperature region that moves inward along with the fragments, while a temperature layer of approximately 100 eV remains at the edge. This inside-out cooling does not occur at higher fragment numbers or lower penetration speed, as discussed later. For pure deuterium injection, due to lack of radiation, the thermal energy is maintained and the temperature profile gradually relaxes over time. In contrast, with neon mixed injections, the core temperature suddenly drops to below 10 eV following the 1/1 MHD activity, and then initiating the inside-out cooling.
3.3 Influence of fragment size
@@
To understand the influence of fragment size on the disruption mitigation effectiveness, simulations with different fragment sizes are conducted with a neon fraction of 0.12% and 10%, corresponding to the cases MF_FV_Ne0.12, SF_FV_Ne0.12, MF_FV_Ne10 and SF_FV_Ne10 from Tab. 1.
@@
Figure 13: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity and (d) radiation power for different fragment numbers for fixed total volume/material with 0.12% neon. The time of fragment arrival $t_{arr}$ and exit $t_{ext}$ at/from the plasma domain are marked by the purple and cyan vertical dashed lines, respectively. The times at the end of TQ are annotated by the vertical dotted lines.
@@
Figure 14: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity and (d) radiation power for different fragment numbers for fixed total volume/material with 10% neon. The time of fragment arrival $t_{arr}$ and exit $t_{ext}$ at/from the plasma domain are marked by the purple and cyan vertical dashed lines, respectively. The times at the end of TQ are annotated by the vertical dotted lines.
@@
Figure 15: Flux-averaged temperature profile evolution for different fragment numbers with 0.12% neon. The dashed lines indicate the estimated trajectory of the bulk fragments.
@@
Figure 16: Flux-averaged temperature profile evolution for different fragment numbers with 10% neon. The dashed lines indicate the estimated trajectory of the bulk fragments.
Figure 17: Electron temperature distribution near the end of the TQ for cases with varying numbers of fragments. As the number of fragments increases, the temperature belt around 100 eV near the edge is progressively eliminated, indicating more effective edge cooling. This enhanced cooling is attributed to the more elongated and sustained plume generated by smaller fragments, which improves assimilation during the initial injection stage.
@@
Figure 13 shows the overview plots for different fragment numbers in the 0.12% neon scenario, where the total injected material is identical across cases. In figure 13 (a), a trend is observed that a larger number of smaller fragments can lead to a shorter TQ duration as indicated by the thermal energy evolution. This agrees with the lower dimensional results we previously obtained between JOREK and INDEX [16]. The larger total surface area for smaller fragments contributes to stronger assimilation until the majority of the fragments exit the plasma at around 3.5 ms. Since the end of ablation is not determined by a full thermal collapse here, but by the leaving of the fragments, the ablation for all cases saturates at almost the same time, which can be clearly seen in figure 13 (c). As the ablation duration is not constrained by the low temperature at TQ, smaller fragments achieve higher assimilation compared to larger ones. Owing to the larger assimilation, the core collapse occurs earlier with smaller fragments and stronger radiation power is observed at the core collapse phase, as shown in figure 13 (d).
@@
Our numerical results deviate from the experimental findings reported by Jachmich [41], where larger fragments with low neon fraction exhibit better material assimilation. Rocket effect and plasmoid drift are promising candidates for explaining this discrepancy. Smaller fragments are more susceptible to the rocket effect due to smaller mass, which accelerates the fragments and deflects their trajectories. When multiple small fragments ablate simultaneously in a localized region, their combined neutral clouds can generate a larger local pressure peaking, and hence stronger plasmoid drift. In experiments [41], smaller fragments tend to have longer pre-TQ duration yet achieve lower assimilation. Simulations using the NGS model indicate that smaller fragments ablate at a higher rate; if their total assimilation remains lower despite longer ablation durations, it is very likely due to the plasmoid drift and the rocket effect expelling the material from the plasma for low neon fraction injections.
@@
Figure 14 shows the same data, but for the 10% neon cases with different fragment numbers. It can be observed that a larger number of smaller fragments leads to quicker assimilation due to the larger total surface area at the beginning, similar to cases with lower neon content. However, a reduced number of larger fragments has a longer ablation duration, resulting in a relatively similar total assimilation level when ablation ceases. In contrast to the 0.12% neon cases, the ablation stops due to the temperature drop during the TQ, which occurs before the fragments leave the plasma on the high field side. By the end of the TQ phase, the 199 fragments case shows the highest assimilation, followed by the 1105 fragments case, with the 53 fragments case having the lowest assimilation. However, the trend is now less pronounced, and the assimilation levels are not significantly different, aligning with experimental observations that the size effect on assimilation is relatively small for pellets with higher neon fractions [41, 20]. This reaffirms our previous hypothesis, since the plasmoid drift is suppressed at high neon fractions. It should be noted that, in those 10% neon cases, a brief decline in assimilation is observed after the TQ. This is due to a transient increase in impurity diffusivity to 4 - 6 times its original value, i.e. $6.4-9.6\ \mathrm{m^{2}/s}$, to overcome numerical difficulties. Once this challenging period passes, the impurity diffusivity is restored to its original value. For the 199 fragments case with 10% neon, the simulation could not proceed beyond the onset of the core collapse due to numerical issues.
@@
The temperature profile evolution for the 0.12% neon cases is shown in figure 15, and for the 10% neon cases in figure 16. The 10% neon case with 199 fragments is excluded from the comparison, as this simulation did not complete the TQ. In the low neon fraction cases, an inside-out cooling pattern is observed across all configurations, owing to limited impurity assimilation at the edge. However, in the high neon fraction cases, as the number of fragments increases, the cooling pattern transitions from inside-out to outside-in. The plume generated by the smaller fragments is more elongated, ensuring a continuous presence of fragments at the edge. Additionally, smaller fragments exhibit more efficient assimilation during the initial stage. Consequently, a cold front moving alongside the pellets inwards is established more effectively with smaller fragments. This effect is clearly illustrated in figure 17, where the temperature belt around 100 eV is progressively diminished as the number of small fragments increases, indicating enhanced edge cooling and a more pronounced cold front.